if this years free cash flow is 500 million and the growth rate of the cash flow is 5% with a required return of 10%, the value of operation is a little over 10 billion. T/F
Answer #1

(Value of Operations = Free cash Flow *(1+Growth)/(Required Rate - Growth) = 500*(1+5%)/(10%-5%) = 10,500 or 10.5 billion

Hence False because value of operations is 10.5 and not 10 billion

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well